

********************************************************************
UARS WEEKLY STATUS REPORT - 30 January 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data during this past week without apparent problems. Also, a
series of SSPP gridded offset maneuvers were performed using the SUSIM
lower spectrometer.
***********
PEM
***********
PEM/AXIS was active from January 18 through January 22. PEM produced
energy deposition profiles from electrons inverted from the AXIS X-ray
measurements taken during this time. PEM/AXIS and PEM/ZEPS (zenith boom
spectrometers and magnetometer) became active on January 29. PEM is
now producing energy deposition profiles from in situ electrons, in situ ions,
and electrons inverted from measured x-rays.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The MLS instrument has been powered off since Dec 27. MLS is scheduled to
begin north viewing operations on Jan 20.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 365 CD-ROM requests as of 30-January-1998.
There was one order received and one order shipped.
THE CDHF loaded DAAC data for HALOE, HRDI, PEM Versions 3 and 4, NMC, and UKMO
created in the month of November 1997.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-854C=525L)
121 days MLS (4.90) PROD (1344R-198C-3F=1146L)
0 day CLAES (7.0) ENG (17R-1F=17L)
9 days WIN (5.10) ENG (17R-17C=0L)
The CDHF will load December 1997 current day DAAC data in February 1998.
The CDHF has disabled accounts that had been determined inactive.
The accounts still continue to forward email, if they have
forwarding set up. The Group Rep can re-establish a desired account.
Pending software upgrades: IDL 5.0, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 09 January 1998 (Orbit 34597, GMT Day 010) through 16
January 1998 (Orbit 34700, GMT Day 016).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 018/0102 34717 AXIS on
MLS 020/0341 34749 On
PEM 022/2332 34791 AXIS off
MLS 022/2332 34791 Off
The beta angle is now increasing toward a maximum angle of 72.6 degrees on
05 February 1998 (DOY 036).
A series of SSPP gridded offset maneuvers requested by SUSIM was performed
on GMT days 017 through 022. The offsets utilized the 5 and 10 arcminute
intervals, as opposed to the 15 and 30 arcminute intervals.
Two Rmanually generatedS Stellar Offset maneuvers requested by SOLSTICE
occurred on GMT days 020 and 022.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.2 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.2 degrees C. Additional battery performance data for
each battery on 17 January 1998 and 23 January 1998 are provided in
Appendix D.
The clock error ranged from -10.0 to +10.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 01
% OUT-OF-LIMITS = 06
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% HRDI 03
% MLS 02
% PEM 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
TRANSPONDER FREQUENCIES:
Transponder A: 98/020 - 2287.497001
Transponder B: 98/014 - 2287.498785
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/022 show the observatory data loss to be
20 hours, 41 minutes, 35.316 seconds (an increase of 01.056 seconds since
last reported). This is a 0.0372 percent data loss which equals a 99.9628
percent data capture for the mission.
The increase of 01.056 seconds occurred on Day 020 and was the result of
hits in the data.
No testing of the new CMS software release (R7B1V2) occurred this week.
A simulation of a reduced size ATC load was performed this week. The test
was conducted by loading the UTTS OBC with a 768-command ATC load. The new
OBC image accepted the reduced ATC load but a large amount of miscompares
(2263) were revealed during the dump comparison process. The first 768
words compared correctly. However, most of the words in locations 768-3071
miscompared apparently because the ground system logic built the ground
reference image as though it expected a 1024-command ATC load. A meeting
with the programmer was conducted on 21 January 1998 in which the test
results were discussed.
Operational processing of level-0 telemetry analysis files continues in
parallel on both the UMPGAL (new MPG Alpha machine) and CDHF computer
systems. No level-0 file transfer failures (between CDHF and UMPGAL)
occurred this week.
The MS434 Rline tapS of the UARS real-time I channel data to the TOMS 511
line appears to be functioning well. The TOMS/UARS re-engineering team
informed UARS that the UARS data looked good on their test system.
The FOT has finalized the details for participating in the IP return-link
telemetry testing. The FOT was informed by MSOCC (Tom Marsh) that the IP
conversion devices were installed into the MSOCC area on 21 January 1998.
The MSOCC DOC operators have been informed of the special procedures which
must be executed to configure the AP to receive IP data from NASCOM. UARS
plans to start viewing the IP data on 27 January 1998.
The tape loading mechanism on the UCMS 4mm tape drive (prime system) broke
on 18 January. The tape drive was repaired on 22 January. Three daily
incremental saves and one MPG save were missed.
SUSIM operations submitted a request for the FOT to schedule six additional
SOLSTICE offsets for the week of 01 February 1998. Both SOLSTICE
operations and the FOT agreed to this request. One offset will take place
each day beginning with GMT Day 031 and ending with GMT Day 036.
A request for another set of series of SSPP gridded offset maneuvers, to
occur on GMT Days 024-029, was also submitted by SUSIM operations.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued. A
significant amount of coding has been completed, only to discover the
redesigned software now far exceeded its memory allocation. Working on
second iteration of redesign to reduce the code size to within the
available memory space.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
A recent projection of the orbit attitude by the Flight Operations Team
(FOT) indicates that an orbit adjust will probably will not be required
until 1999.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 25 February
1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS batteries and their performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
will probably be required in 1999.
FACILITIES
no change
PERSONNEL CHANGES
no change
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 January 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
34694 016/1202 98-015 MPS V/T Level change to 6
34695 016/1332 98-016 MPS V/T Level change to 5
34697 016/1721 98-017 MPS Reset BATSOC to 100
34717 018/0102 98-018 PEM Power ON AXIS
34746 019/2249 97-019 HRDI Heater delay to 200 Amp. mins.
34749 020/0341 98-020 MLS Power ON
34749 020/0432 98-021 HRDI Heater Delay to 25 amp-min
34759 020/2112 98-022 HRDI Heater Delay to 100 amp-min
34784 022/1309 98-024 OBC Clk Adj fr 20.6 to 0.00 msec/day
34791 022/2332 98-025 PEM PEM AXIS Power OFF
34791 022/2332 98-026 MLS Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 January 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
ACTEMPANA2 Y-LO Temp Sensor Electronics 34696-97
HRIFESURVTMP Y-HI HRDI Interferometer Survival Temp 34755,56
HRPSEQBIMON Y-HI HRDI Quiet Bus Current Monitor 34711,20,56
QTSISAIF Y-HI SSPP Baseplate by SISA I/F CONTINUOUS
SOSTMPRNTRN Y-LO SOLSTICE Temperature Entrance 34696-98
WIEUTEMP Y-LO WINDII EU Temperature CONTINUOUS
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
OBSZ1OOL UFLTR Z1 Residual Exceeded 100 Arcsec 34695
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 January 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
34763 021/02:59 (21 Jan 98) from 00.0 to 20.6 msec/day
34784 022/13:18 (22 Jan 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 January 1998
17 Jan 1998 - GMT Day 017 Beta = 4.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.6 0.000 0.0
2 +134.4/ +0.0 +6.11/+5.15 +32.0/-14.4 24.5 1.025 22.6
3 +89.6/ +0.0 +4.83/+3.87 +30.4/-15.2 24.5 1.020 21.9
23 Jan 1998 - GMT Day 023 Beta = 28.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.6 0.000 0.0
2 +89.6/ +0.0 +6.44/+5.15 +32.0/-16.4 24.2 1.022 23.2
3 +67.2/ -5.6 +5.47/+3.87 +31.2/-16.8 24.2 1.024 22.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
(no change)
_______________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 January 1998
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-001 FOT 34729 018/2103 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period